OpenStudy (anonymous):

The lengths of the legs of a right triangle are 15 cm and 20 cm. What is the perimeter of the triangle? Note: You will need to use the Pythagorean theorem to find the hypotenuse. A. 35 cm B. 25 cm C. 60 cm D. 300 cm E. 20 cm

OpenStudy (anonymous):

get to know the 3-4-5 triangle so that you can do the solve this type of triangle. since the legs are a dilation of a scale factor of 5 the hypotenuse must dilate by that scale factor also.. 3*5=15, 4*5=20, 5*5=25. add them up and you have your perimeter.

OpenStudy (aroub):

(15)^2+(20)^2=(h)^2 225+400=h^2 625=h^2 25=h P=15+20+25 P=60

OpenStudy (anonymous):

I GET IT! ;)